OpenStudy (anonymous):

Name the property shown - (8*5)*11=8*(5*11)

OpenStudy (ybarrap):

Options are associative, distributive and commutative, right?

OpenStudy (anonymous):

Yes

OpenStudy (ybarrap):

associative is (a*b)*c=a*(b*c) distributive is a(b+c)=ab + ac commutative is a + b = b + a Which of these does your answer look most like?

OpenStudy (anonymous):

associative. Thank you!

OpenStudy (ybarrap):

np